OpenStudy (jhannybean):

Solve:\[\frac{dy}{dx} =\frac{1+y^2}{1+x^2}\]

OpenStudy (vishweshshrimali5):

1. Take (1+y^2) in LHS and dx in RHS

OpenStudy (anonymous):

It's seperable

OpenStudy (fibonaccichick666):

So, is this a separable eq?

OpenStudy (jhannybean):

Yep.

OpenStudy (vishweshshrimali5):

We get: \[\large{\cfrac{dy}{1+y^2} = \cfrac{dx}{1+x^2}}\] Now integrate both sides

OpenStudy (fibonaccichick666):

Can you find a way to get only x's on one side and only y's on the other?

OpenStudy (jhannybean):

\[\checkmark \checkmark\]

OpenStudy (vishweshshrimali5):

Remember this: \[\large{\int \cfrac{dy}{1+y^2} = \tan^{-1}y + C}\]

OpenStudy (vishweshshrimali5):

Now can you solve this after this ?

OpenStudy (fibonaccichick666):

so, then, ya just integrate wrt the variables on each corresponding side, you got this

OpenStudy (anonymous):

I cant really help you I dont know this yet Im sorry

OpenStudy (jhannybean):

I posted this up for fun :P

OpenStudy (anonymous):

It's symmetric

OpenStudy (vishweshshrimali5):

This method is called variable separation :D

mathslover (mathslover):

Well done, @vishweshshrimali5

OpenStudy (fibonaccichick666):

oh, I thought this was a bit basic for you

OpenStudy (jhannybean):

It's the RHS i'm confused with.

OpenStudy (fibonaccichick666):

what exactly?

OpenStudy (vishweshshrimali5):

Why?

OpenStudy (jhannybean):

\[tan^{-1}(y) = \tan^{-1}(x) + C\] but it's not \(C\) it's written as \(\tan^{-1}(C)\)

OpenStudy (vishweshshrimali5):

See you can write it as you wish

OpenStudy (fibonaccichick666):

That is incorrect notation

mathslover (mathslover):

Constant will remain constant. lol

OpenStudy (anonymous):

thank you for trying to entertain us but I know another way you can entertain me

OpenStudy (fibonaccichick666):

it has to be +C

OpenStudy (vishweshshrimali5):

C is basically just a constant. \[\large{\tan^{-1}(C)}\] would also be a constant

OpenStudy (anonymous):

and that is by teaching me how to do this ;)

OpenStudy (vishweshshrimali5):

You can modify the integration constant as the need arises to make the final equation more presentable :) You can write: C or C^2 or C^3 or sqrt(C) or C/2 and so on

OpenStudy (fibonaccichick666):

But that is not correct, you would have to say \(tan^{-1}(x)+tan^{-1}(C)\) you cannot just say of C alone

OpenStudy (fibonaccichick666):

C can be anything you want, but you must include it in someway

OpenStudy (jhannybean):

The solution here is written as \[\tan^{-1}(y) = \tan^{-1}(x) +\tan^{-1}(C)\]\[y=\tan(\tan^{-1}(x)+\tan^{-1}(C)) = \frac{x+C}{1-Cx}\]

OpenStudy (vishweshshrimali5):

@Jhannybean You can replace C by some other constant say C' And I have chosen: C' = tan^(-1) C

OpenStudy (jhannybean):

Ohh, ok makes sense, makes sense.

OpenStudy (fibonaccichick666):

I agree with the first step, but not sure how they did the simplification to that fraction, there must be a trig sub I am unaware of

OpenStudy (vishweshshrimali5):

It does not matter whether I write C or C' as both are some constants whose value is unknown to me.

OpenStudy (jhannybean):

I was just curious if there was another method in solving this.

OpenStudy (vishweshshrimali5):

As for the simplification: It is just a simple inverse trigonometry formula

OpenStudy (jhannybean):

@Kainui

OpenStudy (fibonaccichick666):

could always use variation of parameters or similar, but the easiest way is the best way here

OpenStudy (vishweshshrimali5):

You can always use substitution

OpenStudy (vishweshshrimali5):

Substitute: \[\large{x = \tan (u)}\] \[\large{\implies u = \tan^{-1} x}\] \[\large{\implies du= \cfrac{1}{1+x^2} dx}\] Similarly: \[\large{y = \tan (v)}\] \[\large{\implies dv = \cfrac{1}{1+y^2}dy}\]

OpenStudy (vishweshshrimali5):

Now your integrals would simplify to: \[\large{\int du = \int dv}\] \[\large{\implies u = v + c}\] Replace them with \(\tan^{-1} x\) and \(\tan^{-1} y\) and you would get the same answer Again you would have to replace c by \(\tan^{-1} c\)

OpenStudy (vishweshshrimali5):

Though, just an important point: I just use the same basic concept in a different way. This method is called substitution method.

OpenStudy (jhannybean):

and why exactly would we have to replace c?

OpenStudy (vishweshshrimali5):

Its just to make your equation more presentable

OpenStudy (fibonaccichick666):

to make it look pretty usually,

OpenStudy (jhannybean):

So it's not a crime to leave it as \(C\).... it would just be considered ambiguous?

OpenStudy (fibonaccichick666):

nope, C is perfectly acceptable

OpenStudy (vishweshshrimali5):

Let me ask you something: Which one do YOU like more ? \[\cfrac{1}{3} x + \cfrac{1}{2} y =1\] or \[2x + 3y = 6\] Similarly: \[\large{x^2 + y^2 = 1}\] or \[\large{r = 1}\]

OpenStudy (jhannybean):

I had understood the rest of the integration process and how to solve this, I was just simply stuck on the constant part, heh.

OpenStudy (jhannybean):

Oh I see @FibonacciChick666 :)

OpenStudy (fibonaccichick666):

Making it pretty and presentable is all there is, now you can use any arbitrary constant but if you do not put it, It WILL be incorrect, so just remember that

OpenStudy (anonymous):

Use trig addition formula for \(\tan\).

OpenStudy (jhannybean):

Yeah ofc.

OpenStudy (fibonaccichick666):

I never heard of a trig addition formula???(or if I did, I totally forgot)

OpenStudy (anonymous):

Sum angle formula

OpenStudy (jhannybean):

oh yeah, that makes sense.\[\tan(\alpha) +\tan(\beta) = \frac{\tan(\alpha) +\tan(\beta)}{1-\tan(\alpha)\tan(\beta)}\]

ganeshie8 (ganeshie8):

\(\tan^{-1}(C_1)\) is as much an arbitrary constant as \(C\)

OpenStudy (kainui):

@vishweshshrimali5 It's been too long since I've seen you! Good to see you're still on top of integrals! @Jhannybean Another example where choosing +C as something else as an example of its utility is in this common integral: \[\Large y= \int\limits \frac{1}{x}dx \\ \Large y= \ln x + C\] if we choose C instead to be ln(C) we have: \[\Large y = \ln x + \ln C = \ln(xC) \\ \Large e^y=Cx\] Which you might be more comfortable rearranging around with completely different variables as \[\Large y=Ce^x\]

OpenStudy (vishweshshrimali5):

Hi @Kainui ! College really makes things difficult in beginning ;) But, now I'm back :D

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @ganeshie8 \(\tan^{-1}(C_1)\) is as much an arbitrary constant as \(C\) \(\color{blue}{\text{End of Quote}}\) Technically it is restricted to a certain image and may have multiple solutions. But where di \(\tan^{-1}(C_1)\) come frome anyway?

OpenStudy (jhannybean):

That's what I was wondering in the first place.

OpenStudy (vishweshshrimali5):

@wio We just replaced an arbitrary constant with another arbitrary constant

OpenStudy (fibonaccichick666):

Though tan^-1(C) isn't arbitrary technically...

OpenStudy (bohotness):

∫1x2dx=2∫1y(y2−3)dy

OpenStudy (fibonaccichick666):

it is restricted

OpenStudy (bohotness):

-1/x+c=ln(y^2-3) -1/1+c=ln(2^2-3) -1+c=ln(1) -1+c=0

ganeshie8 (ganeshie8):

recall the domain of arctan(x) fib

ganeshie8 (ganeshie8):

what restrictions are you guys talking about ? :/

OpenStudy (fibonaccichick666):

domain is, but what we end with isn't then

OpenStudy (vishweshshrimali5):

@FibonacciChick666 arctan(x) is defined for all real values of x

OpenStudy (fibonaccichick666):

em that made no sense, let me say, ok if we just had C we could make it equal to 8 right? but tan^-1{C} can never equal 8

OpenStudy (fibonaccichick666):

we are eliminating answers by using tan^-1 in my opinion

OpenStudy (vishweshshrimali5):

tan^(-1) C can never be 8...uhm. why ?

OpenStudy (fibonaccichick666):

because it is only defined on y=-pi/2, pi/2

ganeshie8 (ganeshie8):

got you point, so are you saying the restriction is going away because we're removing the arctan before reaching final solution ?

OpenStudy (kainui):

@FibonacciChick666 You are right, but if we do that then we would be getting into complex analysis for the full answer, but it's nice to have a simple, real answer to evaluate as well.

OpenStudy (vishweshshrimali5):

What @Kainui said ^^^ :D

OpenStudy (fibonaccichick666):

I am saying that since arctan has a restricted range, we are eliminating, so it would be better and all inclusive just to put C or D or whichever variable you want instead

OpenStudy (jhannybean):

http://www.cliffsnotes.com/assets/11440.jpg

OpenStudy (vishweshshrimali5):

The "function" arctan has the range (-pi/2, pi/2) but that does NOT mean that tan(pi) won't exist...

OpenStudy (anonymous):

If you make an unnecessary restriction on \(C\), then you are eliminate potential solutions. If you make get rid of necessary restrictions for \(C\), then you allow false solutions.

OpenStudy (fibonaccichick666):

^what he said, since the restriction of the range of inv. tan isn't needed, it removes solutions

OpenStudy (vishweshshrimali5):

hmm...

OpenStudy (anonymous):

\[ \tan^{-1}(y)=\tan^{-1}(x)+C\\ y = \tan(\tan^{-1}(x)+C) = \frac{x+\tan(C)}{1-x\tan(C)} = \frac{x+C'}{1-C'x} \]I don't think this is necessarily a bad answer.

OpenStudy (kainui):

We can write \[\Large y = \tan ( \arctan(x) +C)\] However in a physical situation y taking on complex numbers is meaningless.

OpenStudy (fibonaccichick666):

I mean, personally, I would just leave it there^ No need to simplify it. It's done enough

OpenStudy (fibonaccichick666):

and so long as it is +C and not +tan^-1C, I will agree

OpenStudy (kainui):

Yeah, besides indefinite integrals don't really mean anything. In real life you would be solving a problem with initial conditions and such or simply be able to set it to whatever you like just like we arbitrarily assign where 0 potential energy is. It can be at infinity, or ground level, it doesn't matter.

OpenStudy (fibonaccichick666):

but this is math. We never have REAL WORLD problems. NEVER! That's just absurd. **

OpenStudy (jhannybean):

thanks for clearing that up :)

OpenStudy (kainui):

Yeah that's my point, since it's not the real world, let's cut out this ugly answer you wrote and rewrite it with restrictions so it looks all pretty and mathematically elegant. =P

ganeshie8 (ganeshie8):

lol thats so true :P Okay, so aren't we committing the same mistake with C and lnC business or it doesn't matter here because ln(C) produces all real numbers ?

OpenStudy (anonymous):

If you were to just ignore the existence of indefinite integrals all together...\[ \int_C^x \frac{dt}{1+t^2} = \tan^{-1}(x) - \tan^{-1}(C) \]

OpenStudy (kainui):

Yeah, ln(C) is just as good as plain C because they have the same range. For every value that C takes on we have an exact 1 to 1 for every value ln(C) can take on. It just seems weird because the domain of ln(C) is different.

OpenStudy (fibonaccichick666):

agreed